The Parks and Recreation Department of Burkett County estimates that the initial cost of a “barebones” permanent river park will be $2.3 million. Annual upkeep costs are estimated at $120,000. Benefits of $340,000 per year and disbenefits of $40,000 per year have also been identified. Using a discount rate
of 6% per year, calculate (a) the conventional B/C ratio and (b) the modified B/C ratio.
1
Expert's answer
2021-10-06T00:31:21-0400


(1) PW of cost = $1000


PW of benefits"= 300(P/A, 6\%,5) = 300\times4.212 = 1263.6"


B/C ratio = 1.263


(2) PW of cost = $1000


"PW of benefits = 500(P/A, 6\%, 5) - 50(P/G, 6\%,5) = 500x4.212 - 50 x 7.934 = 1709.3"